OpenStudy (anonymous):

Ella's college professors assign homework independently from each other. Ella knows that there is a 60% chance that she will have math homework tonight and a 70% chance that she will have English homework. What is the probability that she will NOT have math or English homework tonight?

ganeshie8 (ganeshie8):

So, she does not want any home work tonight. that means you need to find : P("no math homework" AND "no english homework" )

ganeshie8 (ganeshie8):

since both events are independent, you can simply multiply the probabilities

ganeshie8 (ganeshie8):

P("no math homework" AND "no english homework" ) = P("no math homework") * P("english homework")

OpenStudy (anonymous):

42%?

ganeshie8 (ganeshie8):

nope, try again

OpenStudy (anonymous):

That is what I got when I multiplied.

ganeshie8 (ganeshie8):

you need to find P("no math homework") first

ganeshie8 (ganeshie8):

P("no math homework") = 1- P("getting math homework" = ?

OpenStudy (anonymous):

I need to know how, I don't expect you to give me a hand out, but help?

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

P("no math homework") = 1- P("getting math homework") = 1 - 0.6 = 0.4

ganeshie8 (ganeshie8):

similarly, P("no english homework") = 1- P("getting english homework") = 1 - 0.7 = 0.3

ganeshie8 (ganeshie8):

therefore, P("no math homework" AND "no english homework" ) = 0.4 * 0.3

ganeshie8 (ganeshie8):

that gives u 12%

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense

OpenStudy (anonymous):

That other one I got wrong and then got one just like it, Carson drives to school the same way each day and there are two independent traffic lights on his trip to school. He knows that there is a 30% chance that he will have to stop at the first light and an 80% chance that he will have to stop at the second light. What is the probability that he will NOT have to stop at either light? Do you simply turn them into decimals and then multiply?

OpenStudy (anonymous):

I got 14% on this one.

ganeshie8 (ganeshie8):

probability for "stopping at first light" = 30% that means, probability for "NOT stopping at first light" = \(70\%\)

ganeshie8 (ganeshie8):

probability for "stopping at second light" = 80% that means, probability for "NOT stopping at second light" = \(20\%\)

ganeshie8 (ganeshie8):

So, "probability for NOT stopping at either lights" = 0.7 * 0.2 = 0.14 = \( 14\% \)

ganeshie8 (ganeshie8):

you're right !!

OpenStudy (anonymous):

Thank you so much for your time and actually helping me understand! I appreciate that so much. :)

ganeshie8 (ganeshie8):

np... u wlc :)

OpenStudy (anonymous):

Yeah I got 12% too and it was correct :)